OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find [f(4)] 2.

OpenStudy (anonymous):

lol kinda like that yeah

OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find [f(4)] 2. the twos are squares

OpenStudy (the_fizicx99):

is it like this ? \[[f(4)]^{2} \]

OpenStudy (anonymous):

yes it is

OpenStudy (the_fizicx99):

\[f(x) = x ^{2} + 1 \] and \[g(x) = 3x + 1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then the last two on end of find f..... is a square too

OpenStudy (the_fizicx99):

Ok, See when you have exponents like, \[(x ^{a})^{b}\] than you just multiply them, \[^{a}*^{b} = ^{ab}\] \[(x)^{ab}\] So in, x^2 plug in 4 for x \[f(4) = (4^{2})^{2} \rightarrow 4^{4} \rightarrow 256 \] f(4) = 256 + 1 257 f(4) = 257

OpenStudy (anonymous):

wow so helpful :)

OpenStudy (the_fizicx99):

For g(x) do the same, \[g(4) = 3(4)^{2} + 1 = 12^{2} + 1 \rightarrow 144 + 1 = 145\] g(4) - 145

OpenStudy (the_fizicx99):

g(4) = 145 *